### v1

# Inference on multiple quantiles in regression models by a rank-score approach

Abstract
This paper tackles the challenge of performing multiple quantile regressions across different quantile levels and the associated problem of controlling the familywise error rate, an issue that is generally overlooked in practice. We propose a multivariate extension of the rank-score test and embed it within a closed-testing procedure to efficiently account for multiple testing. Theoretical foundations and simulation studies demonstrate that our method effectively controls the familywise error rate while achieving higher power than traditional corrections, such as Bonferroni.
Keywords — Closed testing; Familywise error rate; Global quantile regression; Rank-score process.
1 Introduction
Quantile regression extends linear regression by estimating a conditional quantile of a response of interest, providing a more comprehensive view of the relationship between covariates and the response beyond just the conditional mean. Fitting multiple quantile regressions at different quantile levels is crucial in many areas where the impact of some covariates varies across the distribution. For example, in finance, investment firms study the effect of macroeconomic variables on the lower quantiles of stock returns during economic downturns, focusing on worst-case scenarios for risk management and loss mitigation. In medical research, the effect of a treatment intended to reduce cholesterol may be more pronounced at the upper quantiles of the cholesterol distribution, particularly for patients with extremely high initial levels. Quantile regressions help identify which patients are most likely to benefit from the treatment, enhancing the effectiveness of medical interventions. In such contexts, analyzing a single quantile value alone could overlook significant results of interest.
We focus on the problem of testing the evidence of a covariate effect across multiple quantile levels. In the literature, this is often referred to as regional quantile testing when the quantile levels form a continuous range, and as global or partial quantile regression when the levels form a finite set of quantiles $\{\tau_{1},\dots,\tau_{K}\}$. Various methods exist for regional quantile testing, with examples found in the work of sun2021model; belloni2019conditional; he2010detection. In contrast, a valid approach for conducting quantile regressions at multiple discrete $K$ levels is lacking. Typically, researchers perform $K$ separate regressions, analyzing each regression’s results and the corresponding $p$-values/test statistics independently.
This practice, however, leads to a multiple testing problem (goeman2014multiple), where performing $K$ separate tests increases the risk of false discoveries if $p$-values are not adjusted to control family errors such as familywise error (FWER) or false discovery rates. The multiplicity problem is not commonly addressed within the quantile regression literature. To the best of our knowledge, only mrkvicka2023globalquantileregression consider multiple testing in this framework; yet, their proposal lacks a solid theoretical foundation.
The strategy of applying quantile regression across various quantiles and making inferences about the effects of covariates without addressing multiple testing issues is a common oversight in various disciplines. Numerous examples can be found in studies spanning markets (wang2012computing), biology (agarwal2019quantile), management (bien2016study; karimi2024r), economics (fousekis2005demand; falk2012quantile), public health (juvanhol2016factors), health (cheah2022malaysian), education (bassett2002quantile), energy (cheng2020quantile), and climate (chamaille2007detecting). This widespread practice highlights a methodological gap, emphasizing the need for the field to develop and apply rigorous statistical methods that properly account for multiple comparisons and ensure the validity of the findings (ioannidis2005most).
We propose a principled yet straightforward framework for performing valid inference across multiple quantile levels. A naive but valid approach would be to apply standard corrections, such as Bonferroni, to the set of $p$-values obtained from the $K$ separate quantile regressions. However, as demonstrated by goeman2021only, only closed testing procedures are admissible for strong FWER control; every other method is either equivalent to, or uniformly dominated by, some closed testing procedure. Consequently, we adopt the closed testing principle as the inferential backbone of our approach. The closed testing framework controls the FWER by testing all possible intersections of the individual hypotheses $H_{0}^{1},\dots,H_{0}^{K}$ using valid local tests. In short, for any $j=1,\dots,K$, we can reject $H_{0}^{j}$ at a pre-determined level $\alpha$ if all intersection hypotheses containing $H_{0}^{j}$ are rejected at that level $\alpha$. The validity and efficiency of the closed testing procedure thus hinge on the choice of these local tests. The commonly used Wald-type test for quantile regression (koenker1996quantile) suffers from finite-sample distortions, especially when heteroscedasticity affects the data, and slow convergence to its asymptotic distribution, even under independent and identically distributed errors (de2025closed; koenker2005quantile).
A robust alternative is the rank-score test (gutenbrunner1993tests; koenker1999goodness; kocherginsky2005practical), whose multivariate version, however, has not yet been fully developed and theoretically investigated. We therefore propose the multivariate extension of the rank-score test, and we integrate it as the local test within the closed testing procedure. This approach increases analytical flexibility and guarantees valid control of type I errors across multiple quantile levels, thereby improving the reliability of statistical inferences from quantile regression analyses.
To facilitate the application of the proposed methodology, we provide an R package named quasar (QUAntile Selective inference Adjustment via Rank-scores) available on CRAN (https://cran.r-project.org/web/packages/quasar/index.html).
The manuscript is organized as follows. Section 2 presents the necessary background on quantile regression, while Section 3 introduces the closed testing procedure. Section 4 details the theory of the multivariate version of the rank-score test. Section 5 presents simulation studies showing that the Wald-type test fails to control the type I error rate, whereas the proposed method effectively controls the FWER while maintaining competitive power compared to traditional corrections such as Bonferroni.
2 Setting
Consider a sample of $n$ independent units. For each $i=1,\dots,n$, we observe a continuous response $y_{ni}\in\mathbb{R}$, a scalar covariate of interest $x_{ni}\in\mathbb{R}$, and a $p$-dimensional vector of nuisance covariates $z_{ni}\in\mathbb{R}^{p}$, where $p=1,\dots,n-2$. The first element of $z_{ni}$ equals $1$ for all units and acts as an intercept. Throughout, we make explicit the dependence on $n$ whenever relevant.
We assume that the following location-shift model (koenker1999goodness) generates the data:
$Y_{ni}|x_{ni},z_{ni}=x_{ni}\beta+z_{ni}^{\top}\gamma+(x_{ni}\theta+z_{ni}^{\top}\eta)U_{ni},$
where $\beta,\theta\in\mathbb{R}$, $\gamma,\eta\in\mathbb{R}^{p}$, and $U_{ni}$ are independent and identically distributed errors from a distribution function $F$. Hence, $Y_{ni}$ is a random variable whose distribution depends on $F$, while $y_{ni}$ represents its realization. Further, denote $\Gamma_{n}$ as a diagonal $n\times n$ matrix with $i$th entry $\sigma_{ni}=|x_{ni}\theta+z_{ni}\eta|$; this specification allows for relaxation over the homoscedasticity assumption of the units. The only requirement we set is $\eta_{1}\neq 0$.
[…]
$Q_{Y_{ni}\mid x_{ni},z_{ni}}(\tau\mid x_{ni},z_{ni})=\beta(\tau)x_{ni}+\gamma(\tau)z_{ni},$
where $\beta(\tau)=\beta+\theta F^{-1}(\tau)$ is the quantile-specific coefficient for ${x}_{ni}$, and $\gamma(\tau)=\gamma+\eta F^{-1}(\tau)$ for ${z}_{ni}$. All results in the manuscript rely on, sometimes implicit, conditioning on the observed covariates that define the design matrix, as is customary in regression-type models.
The purpose of this manuscript is based on the following problem: given a sample of $n$ elements, we want to test the effect of the target covariate at $K$ different quantile levels $\tau_{1},\ldots,\tau_{K}$. Accordingly, we formulate the following set of $K$ null hypotheses, each tested against a two-sided alternative:
$\begin{cases}H_{0}^{1}:\beta(\tau_{1})=0\\ \;\vdots\\ H_{0}^{K}:\beta(\tau_{K})=0.\end{cases}$
For simplicity, we will refer to these null hypotheses simply as hypotheses throughout the paper.
Although testing for the nullity of the coefficients is the most common practice when investigating statistical hypotheses, this manuscript will actually allow for any null values $\beta_{0}(\tau_{1}),\dots,\beta_{0}(\tau_{K})$ of interest. Furthermore, one-sided alternatives can be easily derived, as well.
We introduce some minimal assumptions required for the remainder of the paper. These are not meant to limit practical applicability but to exclude cumbersome or degenerate cases. They are inspired by gutenbrunner1992regression and koenker1999goodness and are standard in inference procedures for quantile regressions.
Assumption 1.
$F$ has continuous density $f$ with respect to the Lebesgue measure, and $f$ is positive and finite on {t:0<F(t)<1t:0<F(t)<1}.
Assumption 2.
The design matrix $M_{n}=\begin{bmatrix}X_{n}&Z_{n}\end{bmatrix}$ of dimension $n\times(p+1)$ contains an intercept.
Assumption 3.
$M_{n}||_{\infty}=o(n^{1/2})$, where $||\cdot||_{\infty}=\max_{j=1,\ldots,p+1}\sum_{i=1}^{n}|m_{ij}|$.
Assumption 4.
$Q_{n}=n^{-1}M^{\top}_{n}M_{n}\xrightarrow{n\xrightarrow{}\infty}Q$, where $Q$ is a $(p+1)\times(p+1)$ matrix.
Assumption 5.
For $n=1,2,\dots$ and $i\leq n$, there exist two finite constants $\lambda_{1},\lambda_{2}>0$ such that $\lambda_{1}<\sigma_{ni}<\lambda_{2}$.
Assumption 6.
$C_{n}=n^{-1}M^{\top}_{n}\Gamma^{-1}_{n}M_{n}\xrightarrow{n\xrightarrow{}\infty}C$, where $C$ is a $(p+1)\times(p+1)$ matrix.
Assumption 7.
$V_{n}=n^{-1}(X_{n}-\hat{X}_{n})^{\top}(X_{n}-\hat{X}_{n})\xrightarrow{n\xrightarrow{}\infty}V$, where $V$ is a scalar and $\hat{X}_{n}=Z_{n}(Z_{n}^{\top}\Gamma_{n}^{-1}Z_{n})^{-1}Z^{\top}_{n}\Gamma^{-1}_{n}X_{n}$.
Assumption 1 formalizes that we restrict our attention to absolutely continuous response variables. Assumption 2 asks that an intercept must always be included in the fitted model. Assumption 3 requires that each entry of the design matrix does not grow too fast; it is weaker than assuming finite elements over the entire design matrix. Assumption 5 posits that the variances of the response vector are bounded away from zero and infinity. Assumptions 4-6-7 guarantee the existence of well-defined limits for some matrices based on the variance-covariance matrix of the covariates.
With these conditions in place, the following section outlines the closed-testing procedure used to achieve strong control of the FWER.
3 Closed testing principle
To test the hypotheses of interest $H_{0}^{1},\ldots,H_{0}^{K}$ with FWER control at a pre-specified level $\alpha\in(0,1)$, we employ the closed testing principle (marcus1976closed). This is the optimal framework for controlling the FWER, as any valid multiple testing method is either equivalent to a closed testing procedure or uniformly improved by one. This optimality property extends beyond FWER control and holds for other error measures, such as the false discovery proportion (goeman2021only).
Let $\mathcal{K}=\{1,\ldots,K\}$ denote the index set of all hypotheses, and let $\mathcal{K}_{0}\subseteq\mathcal{K}$ represent the unknown subset of true hypotheses. The closed testing principle requires considering not only individual hypotheses, but also intersection hypotheses of the form
$H_{0}^{\mathcal{C}}=\bigcap_{j\in\mathcal{C}}H_{0}^{j},\qquad\mathcal{C}\subseteq\mathcal{K}.$
Such an intersection hypothesis is true if and only if all individual hypotheses involved are true, i.e., $\mathcal{C}\subseteq\mathcal{K}_{0}$. Furthermore, the closed testing procedure requires a valid $\alpha$-level local test $\psi:2^{\mathcal{K}}\rightarrow\{0,1\}$, where $2^{\mathcal{K}}$ is the power set of $\mathcal{K}$ and $\psi(\mathcal{C})=1$ denotes rejection of $H_{0}^{\mathcal{C}}$ for $\mathcal{C}\subseteq\mathcal{K}$. The local test is said to be at level $\alpha$ if $\text{pr}(\psi(\mathcal{C})=1)\leq\alpha$ for any $\mathcal{C}\subseteq\mathcal{K}_{0}$, where the probability is taken under the true, unknown data-generating process. By convention, the intersection hypothesis $H_{0}^{\emptyset}$ corresponding to the empty set is always true, and $\psi(\emptyset)=0$.
Based on the local test, closed testing defines a global decision rule
$\bar{\psi}:2^{\mathcal{K}}\rightarrow\{0,1\},\qquad\bar{\psi}=\min\{\psi(\mathcal{E}):\mathcal{C}\subseteq\mathcal{E}\subseteq\mathcal{K}\}.$
Thus, closed testing rejects a hypothesis $H_{0}^{\mathcal{C}}$ whenever the local test rejects all intersection hypotheses corresponding to supersets of $\mathcal{C}$. By construction, the procedure provides strong control of the FWER:
$\text{pr}(\exists\,\mathcal{C}\subseteq\mathcal{K}_{0}:\bar{\psi}(\mathcal{C})=1)\leq\alpha.$
While the procedure applies to all intersection hypotheses, our focus here is, as is typically the case, on testing the individual hypotheses only.
For illustration, consider the following simple example, adapted from goeman2011multiple. Figure 1 shows the tree of all intersection hypotheses for $K=3$, reflecting the subset-superset relationships. Each level corresponds to index sets of the same size, and an arrow from a node $\mathcal{E}$ to a node $\mathcal{C}$ indicates that $\mathcal{C}\subset\mathcal{E}$. The closed testing procedure rejects an individual hypothesis, for example $H_{1}$, only if the local test rejects all intersection hypotheses that include it, namely $H_{\{1\}}=H_{1}$, $H_{\{1,2\}}=H_{1}\cap H_{2}$, $H_{\{1,3\}}=H_{1}\cap H_{3}$, and $H_{\{1,2,3\}}=H_{1}\cap H_{2}\cap H_{3}$.
Figure 1: Tree of intersection hypotheses for three individual hypotheses. Adapted from goeman2011multiple.
In the context of quantile regression, the possible local tests are limited to either tests for individual hypotheses, such as the univariate rank-score test (koenker2005quantile, Section 3.5.4), combined with a Bonferroni correction, or the Wald-type test of koenker2005quantile. However, the Bonferroni correction is known to have notoriously low statistical power (pesarin2001multivariate), while the Wald-type alternative converges slowly to its asymptotic distribution and exhibits distortions in finite samples. These limitations are illustrated, for example, in Section 5 and in de2025closed.
The next section introduces a local test that constitutes the multivariate extension of the test based on regression rank-scores and overcomes the small-sample limitations of the Wald-type approach.
4 Multivariate rank-score test
The proposed local test relies on a rank-based testing approach. For the hypothesis $H_{0}^{j}$ corresponding to a single quantile of interest, the univariate test based on regression rank-scores is described in koenker2005quantile. This is a score-type test, meaning that the test statistic is derived from the sub-model implied by $H_{0}^{j}$. To our knowledge, the literature has focused exclusively on such univariate hypotheses. This section outlines the theoretical proposal of this manuscript, that is, the extension to intersection hypotheses over $k=1,\dots,K$ quantiles, where the case $k=1$ coincides with the cited univariate test. For simplicity of notation, we will refer to the intersection hypothesis corresponding to the first $k$ quantiles $\{\tau_{1},\ldots,\tau_{k}\}$, that is,
$H_{0}^{\{1,\ldots,k\}}=H_{0}^{1}\cap\ldots\cap H_{0}^{k},\qquad H_{0}^{j}:\beta(\tau_{j})=0.$
The first key component of the proposed approach is the vector of regression rank-scores, derived from the estimation procedure of the quantile regression model (koenker2005quantile, Section 6.2) and computed under the sub-model implied by the intersection null hypothesis, to which we will refer as the null model. Let $0<\tau<1$ denote a generic quantile level. The regression rank-scores are
$\hat{a}_{n}(\tau)=\arg\max_{a\in[0,1]^{n}}\{a^{\top}Y_{n}:\;Z^{\top}_{n}a=(1-\tau)Z^{\top}_{n}1_{n}\},$
where $Y_{n}$ denotes the $n$-dimensional vector of the responses and $1_{n}$ is an $n$-dimensional vector of ones. The $n$-vector $\hat{a}_{n}(\tau)$ represents the dual solution to the quantile regression problem
$\hat{\gamma}(\tau)=\arg\min_{\gamma\in\mathbb{R}^{p}}\sum_{i=1}^{n}\rho_{\tau}(y_{ni}-z_{ni}^{\top}\gamma),$
where $\rho_{\tau}(u)=u(\tau-\mathbf{1}\{u<0\})$ denotes the quantile loss function, and $\mathbf{1}\{\cdot\}$ is the indicator function. Notably, the regression rank-scores are $\tau$-dependent as explicitly stated.
As detailed by koenker2005quantile, the score function is obtained by integrating an appropriate score-generating function $\phi:(0,1)\to\mathbb{R}$, which must be non-decreasing and square-integrable. Letting $\hat{a}_{ni}(\tau)$ denote the $i$th component of $\hat{a}_{n}(\tau)$, the score function is thus defined as
$\hat{b}_{ni}(\tau)=\int^{1}_{0}\phi(u)\,d\hat{a}_{ni}(u).$
While gutenbrunner1993tests introduced general score-generating functions $\phi(\cdot)$ for shift-location problems, koenker1999goodness advocated the use of a quantile-specific function for $\tau$-specific problems:
$\phi_{\tau}(t)=\tau-\mathbf{1}\{t<\tau\},$
which yields
$\hat{b}_{ni}(\tau)=\hat{a}_{ni}(\tau)-(1-\tau).$
Let $d_{ni}$ denote the $n$ entries of the vector $D_{n}=X_{n}-\hat{X}_{n}$. Following gutenbrunner1992regression, the quantile rank-score process is
$\hat{W}^{d}_{n}=\left\{\hat{W}_{n}^{d}(\tau)=n^{-1/2}\sum_{i=1}^{n}d_{ni}[\hat{a}_{ni}(\tau)-(1-\tau)],\;0<\tau<1\right\}.$ (1)
This stochastic process forms the basis of the score-based test statistic proposed below. It is characterized by the following theorem from gutenbrunner1992regression.
Theorem 1.
Let Assumptions 1–7 hold and suppose that $H_{0}^{\{1,\ldots,k\}}$ is true. Then
$\hat{W}^{d}_{n}\xrightarrow{d}V^{1/2}W^{*},$
where $W^{*}$ is a Brownian bridge. The convergence in distribution of the process is understood in terms of the Prokhorov topology on the Skorokhod space $\mathbb{D}(0,1)$.
The original theorem is stated for general triangular arrays of possibly multidimensional vectors, with an unspecified structure in place of the elements $d_{ni}$. Theorem 1 presents a more restrictive formulation, sufficient for the purposes of this paper, yet flexible enough to allow extensions, for example, to inference on a multivariate target coefficient.
The distribution of the score vector over multiple quantiles is obtained as follows. Define
$S_{n}=n^{-1/2}(X_{n}-\hat{X}_{n})^{\top}\hat{b}_{n}$
and
$T_{n}=S_{n}^{\top}A^{-1}_{n}S_{n},$
where $A_{n}$ is a $k\times k$ matrix whose $(\ell,r)$th entry equals $V_{n}\delta_{\ell r}$ with $\delta_{\ell r}=\,\min\{\tau_{\ell},\tau_{r}\}-\tau_{\ell}\tau_{r}$.
Theorem 2.
Let Assumptions 1–7 hold and suppose that $H_{0}^{\{1,\ldots,k\}}$ is true. Then $T_{n}\xrightarrow{d}T\sim\chi^{2}_{k}$, where $\chi^{2}_{k}$ denotes a chi-squared distribution with $k$ degrees of freedom.
Proof.
The result follows directly from properties of the Brownian bridge. For two points $\ell,r$, such that $\ell<r$, we have
[…]
where $\mathcal{N}$ denotes the multivariate normal distribution. Multiplying the Brownian bridge by a constant $c$ scales each element of the covariance matrix by $c^{2}$. By Theorem 1 and Slutsky’s theorem, it follows that, for $k$ points in $(0,1)$, the limiting distribution of $S_{n}$ is $k$-variate normal with zero mean vector and covariance matrix $A$, where $A$ is a $k\times k$ matrix whose $(\ell,r)$th entry equals $V\delta_{\ell r}$. The final claim follows from the relationship between the normal and chi-squared distributions. ∎
Theorem 2 enables rank-score-based testing of hypotheses across multiple quantiles simultaneously. Combined with the closed-testing framework described in Section 3, it allows inference on individual quantiles while properly adjusting for multiplicity. Finally, the extension to the multivariate case, where $X_{n}$ has multiple columns, is straightforward, albeit involving more tedious matrix computations.
After establishing the null distribution of the test statistic, we characterize its behavior under alternatives. In particular, its distribution can be derived under local alternatives of the form
$H_{n}^{\{1,\ldots,k\}}=H_{n}^{1}\cap\ldots\cap H_{n}^{k},\qquad H_{n}^{j}:\beta(\tau_{j})={n}^{-1/2}h(\tau_{j}),$
where $h(\cdot)>0$ is a finite constant. In such cases, we obtain the following result.
Theorem 3.
Let Assumptions 1–7 hold and suppose that $H_{n}^{\{1,\ldots,k\}}$ is true. Then $T_{n}\xrightarrow{d}T\sim\chi^{2}_{k}(\zeta)$, where $\chi^{2}_{k}(\zeta)$ denotes a noncentral chi-squared distribution with $k$ degrees of freedom. The noncentrality parameter is $\zeta=g^{\top}A^{-1}g$, where $g\in\mathbb{R}^{k}$ is a vector with components
$g_{j}=C^{11}f\circ F^{-1}(\tau_{j})\,h(\tau_{j}),\quad j=1,\ldots,k.$
Here $f\circ F(\cdot)=f(F(\cdot))$, and $C^{11}=(C^{-1})_{11}$ is the $(1,1)$ element of the inverse of $C$.
Proof.
Let $\hat{W}^{d}_{n}$ be defined as in (1), and
${W}^{d}_{n}=\left\{{W}_{n}^{d}(\tau)=n^{-1/2}\sum_{i=1}^{n}d_{ni}[a^{*}_{ni}(\tau)-(1-\tau)],\;0<\tau<1\right\}$
where $a^{*}_{ni}(\tau)=\mathbf{1}\{U_{ni}>F^{-1}(\tau)\}$.
The proof of Theorem 1 in gutenbrunner1992regression shows that
$\hat{W}_{n}^{d}-W_{n}^{d}=-n^{1/2}\left(\hat{G}^{d}_{n}(\tau)-G^{d}_{n}(\tau)\right)+o_{p}(1),$ (2)
where
$\hat{G}^{d}_{n}(\tau)=n^{-1}\sum_{i=1}^{n}d_{ni}\mathbf{1}\{U_{i}\leq m_{ni}^{\top}\hat{t}_{n}(\tau)/\sigma_{ni}\}$
and
$G^{d}_{n}(\tau)=n^{-1}\sum_{i=1}^{n}d_{ni}\mathbf{1}\{U_{i}\leq m_{ni}^{\top}t(\tau)/\sigma_{ni}\}$
are empirical processes, whose definitions refer to the full model. Here
$\hat{t}_{n}(\tau)=\begin{bmatrix}\hat{\beta}_{n}(\tau)-\beta\\[2.0pt] \hat{\gamma}_{n}(\tau)-\gamma\end{bmatrix},\qquad t(\tau)=\begin{bmatrix}\beta(\tau)-\beta\\[2.0pt] \gamma(\tau)-\gamma\end{bmatrix},\qquad m_{ni}=\begin{bmatrix}x_{ni}\\[2.0pt] z_{ni}\end{bmatrix}$.
Using equation (2) and a first-order linearization with respect to the regression parameters of the difference $\hat{G}^{d}_{n}(\tau)-{G}^{d}_{n}(\tau)$, since $\hat{W}^{d}_{n}$ is evaluated under $H_{0}^{\{1,\ldots,k\}}$ while the true model is given by $H_{n}^{\{1,\ldots,k\}}$, we obtain
$\hat{W}^{d}_{n}-{W}^{d}_{n}=n^{-1/2}f\circ F^{-1}(\tau)D_{n}^{\top}\Gamma_{n}^{-1}Z_{n}\hat{\gamma}_{n}(\tau)-n^{-1/2}f\circ F^{-1}(\tau)D_{n}^{\top}\Gamma_{n}^{-1}M_{n}\begin{bmatrix}h(\tau)\,{n}^{-1/2}\\ {\gamma}(\tau)\end{bmatrix}+o_{p}(1)$
where $M_{n}=\begin{bmatrix}X_{n}&Z_{n}\end{bmatrix}$.
We compare this linearization under the null model and under local alternatives. Under the null model, the relation
$\hat{W}^{d}_{n}-{W}^{d}_{n}=n^{-1/2}f\circ F^{-1}(\tau)D_{n}^{\top}\Gamma_{n}^{-1}Z_{n}\left(\hat{\gamma}_{n}(\tau)-{\gamma}(\tau)\right)+o_{p}(1)$
is used to prove Theorem 1 by gutenbrunner1992regression. On the other hand, under the local alternatives, the additional term
$n^{-1}f\circ F^{-1}(\tau)D_{n}^{\top}\Gamma_{n}^{-1}X_{n}h(\tau)$ (3)
is deterministic. By standard matrix algebra we get
$n^{-1}D_{n}^{\top}\Gamma_{n}^{-1}X_{n}=(C^{-1}_{n})_{11}=C_{n}^{11},$
which is the $(1,1)$ element of the inverse of $C_{n}$.
Hence, using the same argument as in Theorem 2 and Slutsky’s theorem once more, the statistic $S_{n}$ has $k$-variate normal limiting distribution with covariance matrix $A$, as before, and mean vector (3). Further, by standard properties of the normal distribution, the test statistic $T_{n}={S_{n}^{\top}A^{-1}_{n}S_{n}}$ converges in distribution to $T\sim\chi^{2}_{k}(\zeta)$, a noncentral chi-squared distribution, as stated in the theorem. ∎
A few remarks conclude this section. When defining the hypotheses of interest, values different from zero can also be tested. For simplicity, consider a single quantile $\tau_{j}$ and the hypothesis $H_{0}^{j}:\beta(\tau_{j})=\beta_{0}(\tau_{j})$, for a fixed value $\beta_{0}(\tau_{j})\in\mathbb{R}$. The hypothesis is tested by specifying a null model with $\beta(\tau_{j})=\beta_{0}(\tau_{j})$, and optimizing with respect to the remaining regression parameters. This corresponds to the standard use of an offset term in regression modeling. All theoretical results presented in this section remain valid without modification.
[…]
$\hat{f}_{ni}(Q_{ni}(\tau)|z_{ni})=\frac{2h_{n}}{z_{ni}^{\top}\bigl(\hat{\gamma}(\tau+h_{n})-\hat{\gamma}(\tau-h_{n})\bigr)}=\frac{2h_{n}}{d_{ni}}.$ (4)
For the selection of the bandwidth parameter $h_{n}$, we follow the Hall–Sheather rule (hall1988distribution) as adopted by koenker1999goodness. Since the denominator in (4) may not be positive, $\hat{f}_{ni}(Q_{ni}(\tau)|z_{ni})$ is replaced by
$\hat{f}^{+}_{ni}(Q_{ni}(\tau)|z_{ni})=\max\left\{0.01,\frac{2h_{n}}{d_{ni}-\varepsilon}\right\},$
where $\varepsilon>0$ is a small tolerance parameter. This procedure estimates the elements of $\Gamma_{n}$ up to a scale factor independent of $i$, which cancels in the expression for $\hat{X}_{n}$ (koenker1999goodness). The choice of $0.01$ is made for computational reasons and is practically irrelevant.
5 Simulation study
For all simulation studies, we set the sample size to $n=100$ and the significance level to $\alpha=0.05$. For readability, the dependence on $n$ is suppressed in the mathematical notation. Covariate pairs $(x_{i},z_{i})^{\top}$ are drawn from a bivariate normal distribution with mean zero and an equicorrelated covariance matrix with correlation $0.3$, while different generating mechanisms are considered for the responses $y_{i}$. We define five individual hypotheses, each testing whether the effect of $x_{i}$ is zero at the quantile levels $\{0.1,0.25,0.5,0.75,0.9\}$. This results in $2^{5}-1=31$ possible intersection hypotheses, excluding the one corresponding to the empty set, as defined in Section 3. Each simulation scenario is replicated 1,0001{,}000 times.
First, we examine the type I error control of the Wald-type test of koenker2005quantile and the proposed rank-score test. In this setting, $y_{i}\sim\mathcal{N}(0.5\,z_{i},1)$, so there is no effect of $x_{i}$ at any quantile level, and all $31$ intersection hypotheses are tested. The resulting $p$-values, averaged over simulations, are shown in Fig. 2. Only three of the 31 $p$-values fall within the confidence bands for the Wald-type tests, indicating inadequate type I error control, whereas all rank-score $p$-values lie within the bands, showing proper control.
Figure 2: Type I error control for Wald-type and rank-score tests for intersection hypotheses. Average $p$-values; solid grey lines correspond to $\alpha$, and dotted lines represent the $95\%$ simulation confidence interval for the empirical type I error.
Subsequently, we assess and compare the power of the traditional Bonferroni correction and the proposed closed-testing procedure, both applied to rank-score tests. Three data-generating mechanisms are considered:
[…]
(iii) $y_{i}\sim\mathcal{N}\left(0.5+\beta x_{i}+0.5\,z_{i},\,1+|x_{i}|\right).$
Here $t_{5}$ is the Student’s $t$-distribution with $5$ degrees of freedom, while $\mathrm{Exp}(\lambda_{i})$ is the exponential distribution with rate $\lambda_{i}$. The operator $\cdot]_{-30}^{30}$ denotes the value of its argument truncated to the interval [30,30][-30,30]. The parameter of interest, $\beta$, takes values in $\{0.2,0.4,0.6,0.8\}$ in the first two settings, and in $\{0.6,0.8,1.0,1.2\}$ in the third. For any quantile $\tau$, the true coefficient $\beta(\tau)$ depends not only on $\beta$, shared across quantiles, but also on quantile-specific features of the generating distribution. This follows from the definition of the quantile regression coefficient from Section 2.
Figure 3: Empirical power for individual hypotheses. Solid lines represent the closed testing procedure, and dotted lines represent the Bonferroni correction, both applied to rank-score tests.
The power behavior, shown in Fig. 3, is evaluated for each quantile and computed as the proportion of simulations in which an effect of $x_{i}$ is detected. The three scenarios reveal several common patterns. Overall, our procedure attains higher power than Bonferroni, with only minor exceptions, and ensures consistent convergence toward the maximum attainable power as the true parameter increases. When examining individual quantiles, power gains are most pronounced for the more challenging ones, corresponding to regions of the distribution with lower density. This pattern emerges in both tails for the first and third symmetric scenarios, and in the right tail for the second scenario, due to the skewness of the exponential distribution. Analogous results, not reported to avoid repetitions, were obtained for $\gamma=0$ and for correlations between $x_{i}$ and $z_{i}$ equal to 0 and $0.6$, indicating that the findings are robust to changes in nuisance parameters and covariate dependence.
Supplementary material
The code used in the simulation study is available at https://github.com/angeella/quasar/tree/main/simulations.